# Circle bundles with PSC over large manifolds

### v3

@@ 2.3. Donaldson divisors of $M\times S_{g}$
The map $f:Z\to M$, obtained by the composition
[…]
@@
Take any regular value $p\in M$. Then the oriented count #$\#f^{-1}(p)$ is equal to the oriented intersection number #$\#([Z]\cap[\{p\}\times S_{g}])$. Therefore, we have
$\deg(f)$ =#$=\#([Z]\cap[\{p\}\times S_{g}])$
@@
The last inequality holds because $\{p\}\times S_{g}$ is a symplectic submanifold of $(M\times S_{g},\omega)$. ∎
@@ 2.4. Symplectically aspherical manifolds
This class of manifolds was originally introduced by Floer in the context of the Arnol’d conjecture in symplectic geometry. We refer the reader to the survery article [KRT08] for detailed information. The key property of symplectically aspherical manifolds that we shall be using is the following:
@@
Let $(X^{2n},\omega)$ be an integral symplectic manifold which is symplectically aspherical. Then $X$ is a rationally essential manifold. [^3] A $d$-dimensional manifold $M^{d}$ is rationally essential if for the map $\phi:M\to K(\pi_{1}(M),1)$ classifying the universal cover of $M$, $\phi_{*}[M]\neq 0\in H_{d}(K(\pi_{1}(M),1);\mathbf{Q})$
@@
In this way, we obtain a line bundle $L^{\prime}$ over $K(\pi_{1}(X),1)$ such that $i^{*}L^{\prime}=L$. Let us denote $\alpha:=c_{1}(L^{\prime})\in H^{2}(K(\pi_{1},1);\mathbf{Z})$. Then, $i^{*}\alpha=[\omega]$. Consequently, $i^{*}\alpha^{\wedge n}=[\omega^{\wedge n}]\neq 0$. Thus, $i^{*}:H^{2n}(K(\pi_{1}(X),1);\mathbf{Z})\to H^{2n}(X;\mathbf{Z})$ is nonzero. Since $[\omega^{\wedge n}]$ is not torsion, $\alpha^{\wedge n}$ must not be torsion either. Therefore, $i_{*}:H^{2n}(K(\pi_{1},1);\mathbf{Q})\to H^{2n}(X;\mathbf{Q})$ is nonzero. Therefore, by the universal coefficients theorem, $i_{*}:H_{2n}(X;\mathbf{Q})\to H_{2n}(K(\pi_{1}(X),1);\mathbf{Q})$ must be nonzero as well. ∎
@@ 3.1. Proof of Theorem A
Let $W^{n}$ be a compact $n$-dimensional manifold with boundary $E=\partial W$. Suppose
@@
    The inclusion $i:E=\partial W\hookrightarrow W$ induces an isomorphism in $\pi_{1}$.
Then, $\dim_{mc}\widetilde{E}\leq k$.
@@
Let $r:W\to K$ denote the deformation retract. Let $f=r\circ i:E\to K$ denote the composition with the inclusion $i:E\hookrightarrow W$. Since $i$ and $r$ both induce an isomorphism in $\pi_{1}$, so does $f$. Therefore, $f$ lifts to a map between the respective universal covers:
$\widetilde{f}:\widetilde{E}\to\widetilde{K}$
Choose an arbitrary Riemannian metric $g$ on $W$, and let $\widetilde{g}$ be its lift to the universal cover $\widetilde{W}$. We will show that the fibers of $\widetilde{f}$ are uniformly bounded with respect to the metric $\widetilde{g}$. This will give us the desired conclusion, i.e. $\dim_{mc}\widetilde{E}\leq k$.
To this end, let us assume without loss of generality that $f$ is a simplicial map. This can be arranged by the simplicial approximation theorem. Let $\widetilde{p}\in\widetilde{K}$ be a point lying over $p\in K$. Then, we claim $\widetilde{f}^{-1}(\widetilde{p})$ is isometric to $f^{-1}(p)$. Indeed, note that $f^{-1}(p)\hookrightarrow W$ induces the zero map on $\pi_{1}$, since otherwise there would be a loop $\gamma\subset f^{-1}(p)$ which is not nullhomotopic in $W$ such that $f_{*}[\gamma]=0\in\pi_{1}(K)$. This would be absurd, as $f$ induces an isomorphism on $\pi_{1}$. Thus, $f^{-1}(p)$ lifts homeomorphically in the cover to a unique subset in $\widetilde{Y}$ containing $\widetilde{p}$. By uniqueness, this lift is exactly $\widetilde{f}^{-1}(\widetilde{p})$. Consequently, the covering projection must be a homeomorphism:
$\widetilde{f}^{-1}(\widetilde{p})\to f^{-1}(p)$
Since the metric on the cover is $\widetilde{g}$, obtained by lifting the metric $g$ on $W$, the aforementioned homeomorphism is the desired isometry. Thus, the fibers of $\widetilde{f}$ are isometric to the fibers of $f$, which are uniformly bounded as $E$ is compact. Therefore, $\dim_{mc}\widetilde{E}\leq k$. ∎
@@
Let $(M^{2n+2},\omega)$ be an integral symplectic manifold. Let $L$ be the pre-quantum line bundle on $(M,\omega)$. Let $Z^{2n}\subset(M,\omega)$ be a Donaldson divisor, with $[Z]=\mathrm{PD}(k\cdot[\omega])$, for $k\gg 1$ sufficiently large. We set $E_{Z}:=\mathbb{S}(L^{\otimes k}|_{Z})$ as the unit circle bundle of $L^{\otimes k}$ restricted to $Z$ (as introduced in Lemma 2.1).
@@
    Let $\nu(Z)$ be a tubular neighborhood of $Z$ in $X$. Consider the manifold with boundary $W:=M\setminus\nu(Z)$. From Lemma 2.2, $W$ has a handle decomposition consisting of handles of index $\leq n$. Let $V:=W\setminus B^{2n}$ be the $2n$-manifold obtained from removing a $2n$-ball $B^{2n}\subset W^{\circ}$ from the interior of $W$. Then $V$ is a cobordism between $S^{2n-1}$ and $\partial\nu(Z)$. But $\partial\nu(Z)=S(L^{\otimes k}|_{Z})$ by Lemma 2.1. Note that $V$ has handles of index at most $n$. Therefore, $\mathbb{S}(L^{\otimes k}|_{Z})$ is obtained from $S^{2n-1}$ by surgeries of codimension $\geq n\geq 3$. Since $S^{2n-1}$ is PSC, so is $\mathbb{S}(L^{\otimes k}|_{Z})$ by the Gromov-Lawson-Schoen-Yau surgery theorem [SY79, GL80] which states that any manifold obtained from a PSC manifold by surgeries of codimesion $\geq 3$ is also PSC.
@@
    Once again, let us consider the manifold with boundary $W:=M\setminus\nu(Z)$ where $\nu(Z)$ denotes a tubular neighborhood of $Z$ in $X$. Note that $\partial W=\partial\nu(Z)$. By Lemma 2.1, $\partial\nu(Z)\cong\mathbb{S}(L^{\otimes k}|_{Z})=E_{Z}$. Therefore $\partial W\cong E_{Z}$. In Corollary 3, we deduced that $(W^{2n+2},E_{Z})$ satisfies the hypothesis of Lemma 3.1 with $k=n+1$. Therefore, we conclude $\dim_{mc}(\widetilde{E_{Z}})\leq n+1$.
@@
    Since $(M,\omega)$ is symplectically aspherical and $Z\subset M$ is a symplectic submanifold, $Z$ is also symplectically aspherical (see, Example 6). Therefore, by Proposition 7, $Z$ is a rationally essential manifold. By Lemma 2, $\pi_{1}(Z)\cong\pi_{1}(M)$. By hypothesis, $\pi_{1}(M)$ is amenable. Therefore, $Z^{2n}$ is a rationally essential manifold with amenable fundamental group. Hence, by [Dra11, Theorem 7.2], $\dim_{mc}Z=2n$.
@@
    Here $M=N\times S_{g}$, with symplectic form $\omega=\omega_{N}\oplus\omega_{0}$, for some integral symplectic manifold $(N,\omega_{N})$ and a positive integral area form $\omega_{0}$ on $S_{g}$. Choose a Donaldson divisor $Z\subset(N\times\Sigma_{g},\omega)$ such that $[Z]=\mathrm{PD}(k\cdot[\omega])$, for some $k\gg 1$ sufficiently such that, moreover, $k$ is even. Let $f:Z\to N$ be the restriction of the projection map $M=N\times S_{g}\to N$ to $Z$. By Proposition 5 the map $f:Z\to N$ is of positive degree. By hypothesis, $N$ is enlargeable. Therefore, to show that $Z$ is also enlargeable, we just need to show that there is a finite cover $\widehat{Z}\to Z$ such that $\widehat{Z}$ is spin.
@@
    Let $\widehat{\omega}:=p^{*}\omega$ denote the lift of the symplectic form on $M$ to $\widehat{M}$. Then, $(\widehat{M},\widehat{\omega})$ is a symplectic manifold, and $\widehat{j}:\widehat{Z}\to\widehat{M}$ is a symplectic embedding. From (2.3) in Proposition 4, we have:
@@
    $c_{1}(\widehat{Z})=j^{*}c_{1}(\widehat{M})-k\cdot\widehat{j}^{*}[\widehat{\omega}]$ (3.2)
@@
Finally, it follows from Proposition 4 that for distinct, suffficiently large integers $k\neq l$, the corresponding Donaldson divisors $Z_{k},Z_{l}\subset(M,\omega)$ are not homeomorphic. Thus, one finds infinitely many manifolds $Z^{2n}$ satisfying Conclusions $(1)$-$(4)$. ∎
@@ 3.2. Proof of Theorem B
The proof of this result was communicated to us by Hanke [Han25]. We begin by stating the following proposition.
@@
Let $M^{3}$ be a closed orientable $3$-manifold such that $M$ is either enlargeable or $\dim_{mc}(\widetilde{M})=3$. Then the Kneser-Milnor prime decomposition of $M$ must contain an aspherical summand.
@@
Next, we show $\dim_{mc}(\widetilde{M})\neq 3$. Consider the homomorphism
@@
Next, observe that the universal cover of $M$ agrees with the universal cover of $\widehat{M}$. However, the universal cover of $\widehat{M}\cong\#^{m}(S^{1}\times S^{2})$ is obtained from the standard Cayley graph of $F_{m}$ by assigning vertex spaces as $S^{3}$, edge spaces as $S^{2}\times I$, and tubing the two vertex spaces corresponding to the endpoints of any given edge by an $S^{2}\times I$. Consider the map $f:\widetilde{M}\to Cay(F_{k})$ obtained by collapsing down the vertex spaces to the vertices and edge spaces to the edges of the Cayley graph. Thus, the fibers of $f$ have uniformly bounded diameter. Therefore, $\dim_{mc}(\widetilde{M})\leq 1$. In particular, $\dim_{mc}(\widetilde{M})\neq 3$, as desired. ∎
We now proceed to the proof of Theorem B:
@@
Let $M^{3}$ be a closed $3$-manifold such that $M$ is either enlargeable or $\dim_{mc}(\widetilde{M})=3$, and let $E\to M$ be a circle bundle over $M$ such that the total space of $E$ admits a PSC metric. As a first step, we argue that it can be ensured without loss of generality that $(1)$ $M$ is orientable, $(2)$ the circle bundle $E\to M$ is fiberwise orientable.
@@
    If $M$ is enlargeable, then $M$ is orientable by definition. If $\dim_{mc}(\widetilde{M})=3$ and $M$ is not orientable, we pass to the orientation double cover of $M$ and pull the bundle $E\to M$ back to this cover. As the macroscopic dimension of a manifold is a quantity that only depends on the universal cover of the manifold, the cover continues to have macroscopic dimension $3$. Moreover, the total space of the pullback of $E\to M$ is a double cover of the total space of $E$. Therefore, the total space of the pulled back bundle also admits a PSC metric. Therefore, without loss of generality, we may assume $M$ is orientable.

### v5

@@ 2.3. Donaldson divisors of ${N}\times S_{g}$
The map $f:Z\to{N}$, obtained by the composition
$Z\hookrightarrow{N}\times S_{g}\xrightarrow{\pi}{N},$
@@
Take any regular value $p\in{N}$. Then the oriented count $f^{-1}(p)$ is equal to the oriented intersection number $([Z]\cap[\{p\}\times S_{g}])$. Therefore, we have
$\deg(f)$ $=\#([Z]\cap[\{p\}\times S_{g}])$
@@
The last inequality holds because $\{p\}\times S_{g}$ is a symplectic submanifold of $({N}\times S_{g},\omega)$. ∎
@@ 2.4. Symplectically aspherical manifolds
This class of manifolds was originally introduced by Floer in the context of the Arnol’d conjecture in symplectic geometry. We refer the reader to the survey article [KRT08] for detailed information. The key property of symplectically aspherical manifolds that we shall be using is the following:
@@
Let $(X^{2n},\omega)$ be a closed integral symplectic manifold which is symplectically aspherical. Then $X$ is a rationally essential manifold. [^5] A $d$-dimensional manifold $M^{d}$ is rationally essential if for the map $\phi:M\to K(\pi_{1}(M),1)$ classifying the universal cover of $M$, $\phi_{*}[M]\neq 0\in H_{d}(K(\pi_{1}(M),1);\mathbf{Q})$
@@
In this way, we obtain a line bundle $L^{\prime}$ over $K(\pi_{1}(X),1)$ such that $i^{*}L^{\prime}=L$. Let us denote $\alpha:=c_{1}(L^{\prime})\in H^{2}(K(\pi_{1},1);\mathbf{Z})$. Then, $i^{*}\alpha=[\omega]$. Consequently, $i^{*}{\alpha^{\smile n}}=[\omega^{\wedge n}]\neq 0$. Thus, $i^{*}:H^{2n}(K(\pi_{1}(X),1);\mathbf{Z})\to H^{2n}(X;\mathbf{Z})$ is nonzero. Since $\omega^{\wedge n}]$ is not torsion, $\alpha^{\smile n}$} must not be torsion either. Therefore, $i_{*}:H^{2n}(K(\pi_{1},1);\mathbf{Q})\to H^{2n}(X;\mathbf{Q})$ is nonzero. Therefore, by the universal coefficients theorem, $i_{*}:H_{2n}(X;\mathbf{Q})\to H_{2n}(K(\pi_{1}(X),1);\mathbf{Q})$ must be nonzero as well. ∎
@@ 3.1. Proof of Theorem A
Let $W^{{m}}$ be a compact m{m}-dimensional manifold with boundary $E=\partial W$. Suppose
@@
    The inclusion $i:E=\partial W\hookrightarrow W$ induces an isomorphism on $\pi_{1}$.
Then, $\dim_{\mathrm{MC}}(\widetilde{E\phantom{.}}\!)\leq k$.
@@
Let $r:W\to K$ denote the deformation retract. Let $f=r\circ i:E\to K$ denote the composition with the inclusion $i:E\hookrightarrow W$. Since $i$ and $r$ both induce an isomorphism on $\pi_{1}$, so does $f$. Let $q_{E}:\widetilde{E\phantom{.}}\!\to E$ and $q_{K}:\widetilde{K}\to K$ denote the universal covers of $E$ and $K$ respectively. Since $f$ induces an isomorphism on the fundamental group, it lifts to a map between the respective universal covers:
$\widetilde{f}:\widetilde{E\phantom{.}}\!\to\widetilde{K}$,
such that $q_{K}\circ\widetilde{f}=f\circ q_{E}$. Choose an arbitrary Riemannian metric $g$ on E{E}, and let $\widetilde{g}$ be its lift to the universal cover $\widetilde{E\phantom{.}}$}\!. We will show that the fibers of $\widetilde{f}$ are uniformly bounded.
By the simplicial approximation theorem, we may assume without loss of generality that $f$ is a simplicial map with respect to some triangulation of $E$ and a subdivision of $K$. Note that $K$ is compact as it is a retract of a compact manifold $W$. Therefore, $K$ contains finitely many vertices $v_{\alpha}\in K^{(0)}$, $\alpha\in I$ where $I$ is a finite set of indices. Let $B_{\alpha}\subset K$ denote the open star of $v_{\alpha}$. By passing to an even finer subdivision of $K$ if necessary, we may assume without loss of generality that $B_{\alpha}$ is contractible for all $\alpha\in I$. Let $U_{\alpha}:=f^{-1}(B_{\alpha})\subset E$.
We claim that the universal cover $q_{E}:\widetilde{E\phantom{.}}\!\to E$ trivializes over $U_{\alpha}$ for all $\alpha\in I$. It suffices to show that the inclusion $U_{\alpha}\hookrightarrow E$ induces the zero map on the fundamental group. To this end, let $\gamma\subset U_{\alpha}$ be a loop. Then $f_{*}[\gamma]=0\in\pi_{1}(B_{\alpha})$ since $B_{\alpha}$ is contractible. As $f$ induces an isomorphism on $\pi_{1}$, $\gamma]=0\in\pi_{1}(E)$. This proves the claim.
Let $\widetilde{K}$ be equipped with the simplicial structure obtained from lifting the simplicial structure on $K$. For each $\alpha\in I$, we denote the vertices of $\widetilde{K}$ contained in the fiber $q_{K}^{-1}(v_{\alpha})$ as $w_{\alpha,\beta}$, indexed by $\beta\in J$. Then
$q_{E}^{-1}(U_{\alpha})=\bigsqcup_{\beta\in J}V_{\alpha,\beta},$
where $f(V_{\alpha,\beta})\subset\widetilde{K}$ contains $w_{\alpha,\beta}$ and the restriction $q_{E}:V_{\alpha,\beta}\to U_{\alpha}$ is a diffeomorphism. In particular, $q_{E}:(V_{\alpha,\beta},\widetilde{g})\to(U_{\alpha},g)$ is a Riemannian isometry for all $\alpha\in I,\beta\in J$. Therefore, the intrinsic diameter of $V_{\alpha,\beta}$ is equal to the intrinsic diameter of $U_{\alpha}$ for all $\alpha\in I,\beta\in J$. Let $C>0$ denote the maximum of the intrinsic diameters of $(U_{\alpha},g)$, $\alpha\in I$. This exists, as $I$ is a finite set. Since the extrinsic diameter is at most the intrinsic diameter for any open subset in a Riemannian manifold, we obtain $\mathrm{diam}(V_{\alpha,\beta})\leq C$ for all $\alpha\in I,\beta\in J$.
Let $x\in\widetilde{K}$. Since $\{B_{\alpha}\}$ is an open cover of $K$, $q_{K}(x)\in B_{\alpha}$ for some $\alpha\in I$. Therefore,
$\widetilde{f}^{-1}(x)\subset\widetilde{f}^{-1}(q_{K}^{-1}(B_{\alpha}))=q_{E}^{-1}(f^{-1}(B_{\alpha}))=q_{E}^{-1}(U_{\alpha})$
Thus, there exists $\beta\in J$ such that $\widetilde{f}^{-1}(x)\subset V_{\alpha,\beta}$. Hence, $\mathrm{diam}(\widetilde{f}^{-1}(x))\leq\mathrm{diam}(V_{\alpha,\beta})\leq C$. Note that $C$ only depends on the simplicial structure on $K$, the map $f$ and the Riemannian manifold $(E,g)$. Thus, the fibers of $\widetilde{f}$ are uniformly bounded. Furthermore, since $f$ has compact domain and target, and is simplicial, it is $\lambda$-Lipschitz for some $\lambda>0$. Thus, $\widetilde{f}$ is also $\lambda$-Lipschitz. Therefore, $\dim_{\mathrm{MC}}(\widetilde{E\phantom{.}}\!)\leq k$.∎
@@
Let $(M^{2n+2},\omega)$ be a closed integral symplectic manifold. Let $L$ be the pre-quantum line bundle on $(M,\omega)$. Let $Z^{2n}\subset(M,\omega)$ be a Donaldson divisor, with $Z]=\mathrm{PD}(k\cdot[\omega])$, for $k\gg 1$ sufficiently large. We set $E_{Z}:=\mathbb{S}(L^{\otimes k}|_{Z})$ as the unit circle bundle of $L^{\otimes k}$ restricted to $Z$ (as introduced in Lemma 2.1).
@@
    Let $\nu(Z)$ be a tubular neighborhood of $Z$ in $X$. Consider the manifold with boundary $W:=M\setminus\nu(Z)$. From Lemma 2.2, $W$ has a handle decomposition consisting of handles of index $\leq n+1$. Let $V:=W\setminus{B^{2n+2}}$ be the $2n$-manifold obtained from removing a $(2n+2)$-ball $B^{2n+2}\subset W^{\circ}$ from the interior of $W$. Then $V$ is a cobordism between $S^{2n+1}$ and $\partial\nu(Z)$. But $\partial\nu(Z)=\mathbb{S}(L^{\otimes k}|_{Z})$} by Lemma 2.1. Note that $V$ has handles of index at most $n+1$. Therefore, $\mathbb{S}(L^{\otimes k}|_{Z})$ is obtained from $S^{2n+1}$} by surgeries of codimension $\geq{n+1}\geq 3$. Since $S^{2n+1}$} is PSC, so is $\mathbb{S}(L^{\otimes k}|_{Z})$ by the Gromov-Lawson-Schoen-Yau surgery theorem [SY79, GL80] which states that any manifold obtained from a PSC manifold by surgeries of codimension $\geq 3$ is also PSC.
@@
    Once again, let us consider the manifold with boundary $W:=M\setminus\nu(Z)$ where $\nu(Z)$ denotes a tubular neighborhood of $Z$ in $X$. Note that $\partial W=\partial\nu(Z)$. By Lemma 2.1, $\partial\nu(Z)\cong\mathbb{S}(L^{\otimes k}|_{Z})=E_{Z}$. Therefore $\partial W\cong E_{Z}$. In Corollary 3, we deduced that $(W^{2n+2},E_{Z})$ satisfies the hypothesis of Lemma 3.1 with $m=2n+2$ and $k=n+1$. Therefore, we conclude $\dim_{\mathrm{MC}}(\widetilde{E_{Z}})\leq n+1$.
@@
    Since $(M,\omega)$ is symplectically aspherical and $Z\subset M$ is a symplectic submanifold, $Z$ is also symplectically aspherical (see, Example 6). Therefore, by Proposition 7, $Z$ is a rationally essential manifold. By Proposition 2, $\pi_{1}(Z)\cong\pi_{1}(M)$. By hypothesis, $\pi_{1}(M)$ is amenable. Therefore, $Z^{2n}$ is a rationally essential manifold with amenable fundamental group. Hence, by [Dra11, Theorem 1.6], $\dim_{\mathrm{MC}}(Z)=2n$.
@@
    Here $M=N\times S_{g}$, with symplectic form $\omega=\omega_{N}\oplus\omega_{0}$, for some closed integral symplectic manifold $(N,\omega_{N})$ and a positive integral area form $\omega_{0}$ on $S_{g}$. Choose a Donaldson divisor $Z\subset(N\times\Sigma_{g},\omega)$ such that $Z]=\mathrm{PD}(k\cdot[\omega])$, for some $k\gg 1$ sufficiently large such that, moreover, $k$ is even. Let $f:Z\to N$ be the restriction of the projection map $M=N\times S_{g}\to N$ to $Z$. By Proposition 5 the map $f:Z\to N$ is of positive degree. By the Whitney approximation theorem, we may homotope $f$ slightly to be a smooth map without altering the degree. Since $Z$ is compact, the resulting map is $C$-Lipschitz for some $C>0$. Therefore, to show that $Z$ is also enlargeable, we just need to show that there is a finite cover $\widehat{Z}\to Z$ such that $\widehat{Z}$ is spin.
@@
[…]
@@
    $c_{1}(\widehat{Z})={\widehat{j}^{*}c_{1}(\widehat{M})}-k\cdot\widehat{j}^{*}[\widehat{\omega}$] (3.2)
@@
Finally, it follows from Proposition 4 that for distinct, sufficiently large integers $k\neq l$, the corresponding Donaldson divisors $Z_{k},Z_{l}\subset(M,\omega)$ are not homeomorphic. Thus, one finds infinitely many manifolds $Z^{2n}$ satisfying Conclusions (1)(1)-$(4)$. ∎
@@ 3.2. Proof of Theorem B
We begin by stating the following proposition. The proof of this result was communicated to us by Hanke [Han25].
@@
Let $M^{3}$ be a closed orientable $3$-manifold such that $M$ is either enlargeable or $\dim_{\mathrm{MC}}(\widetilde{M})=3$. Then the Kneser-Milnor prime decomposition of $M$ must contain an aspherical summand.
@@
Next, we show $\dim_{\mathrm{MC}}(\widetilde{M})\neq 3$. Consider the homomorphism
@@
Next, observe that the universal cover of $M$ agrees with the universal cover of $\widehat{M}$. However, the universal cover of $\widehat{M}\cong\#^{m}(S^{1}\times S^{2})$ is obtained from the standard Cayley graph of $F_{m}$ by assigning vertex spaces as $S^{3}$, edge spaces as $S^{2}\times I$, and tubing the two vertex spaces corresponding to the endpoints of any given edge by an $S^{2}\times I$. Consider the map $f:\widetilde{M}\to Cay(F_{k})$ obtained by collapsing down the vertex spaces to the vertices and edge spaces to the edges of the Cayley graph. Thus, $f$ is Lipschitz and the fibers of $f$ have uniformly bounded diameter. Therefore, $\dim_{\mathrm{MC}}(\widetilde{M})\leq 1$. In particular, $\dim_{\mathrm{MC}}(\widetilde{M})\neq 3$, as desired. ∎
We record a lemma that will be used in the proof of Theorem B. The proof of the lemma follows directly from the proof of [Bol06, Theorem 2.1]. We provide the details below for completeness. For background on characteristic classes, we refer the reader to [MS74]. See also the proof of [Bol06, Theorem 2.1] which highlights the key notions and properties used. We adopt the following convention: given a class $\alpha\in H^{i}(X)$, we denote $|\alpha|=i$.
Lemma 3.2.
Let $M$ be an orientable $3$-manifold such that its Kneser-Milnor prime decomposition contains an aspherical summand. Let $p:E\to M$ be a fiberwise orientable circle bundle. Suppose the fibers of $p$ are $\pi_{1}$-null in the total space of $E$. Then the classifying map $f_{E}:E\to K(\pi_{1}(E),1)$ for the universal cover of $E$, can not be homotoped into the $2$-skeleton of $K(\pi_{1}(E),1)$.
Proof.
Applying the homotopy fiber long exact sequence to $S^{1}\hookrightarrow E\stackrel{{\scriptstyle p}}{{\to}}M$, we obtain
$\pi_{2}(M)\to\pi_{1}(S^{1})\to\pi_{1}(E)\stackrel{{\scriptstyle p_{*}}}{{\to}}\pi_{1}(M)\to 0$
Since $\pi_{1}(S^{1})\to\pi_{1}(E)$ is zero by hypothesis, $p$ must induce an isomorphism on $\pi_{1}$. Let $q:D(E)\to M$ denote the $D^{2}$-bundle over $M$ associated to the circle bundle $p:E\to M$, so that $E=\partial D(E)$ and $q|_{E}=p$. Since $q$ is a homotopy equivalence, the inclusion $E=\partial D(E)\hookrightarrow D(E)$ induces an isomorphism on $\pi_{1}$. Thus, let us denote
$G:=\pi_{1}(E)\cong\pi_{1}(D(E))\cong\pi_{1}(M)$
Let $f_{M}:M\to K(G,1)$ and $f_{D(E)}:D(E)\to K(G,1)$ denote the classifying maps for the universal covers of $M$ and $D(E)$, respectively. Then $f_{D(E)}$ restricted to $E=\partial D(E)\subset D(E)$ is homotopic to $f_{E}$, and $f_{D(E)}$ restricted to the zero section $M\subset D(E)$ is homotopic to $f_{M}$.
From the hypothesis, it follows that M=N#KM=N\#K for an orientable aspherical $3$-manifold $K$. Let $\varphi:K(G,1)\to K(\pi_{1}(K),1)\simeq K$ be induced from the quotient map to the free factor $G\cong\pi_{1}(N)*\pi_{1}(K)\to\pi_{1}(K)$. Let $\psi:K\to S^{3}$ be a degree $1$ map, given by collapsing the complement of a ball in $K$. Then, let $g:K(G,1)\to S^{3}$ by defined by $g:=\psi\circ\varphi$.
If $f_{E}$ was homotopic into the $2$-skeleton of $K(G,1)$, then $g\circ f_{E}$ would be null-homotopic. Therefore, $g\circ f_{D(E)}$ restricted to $E=\partial D(E)\subset D(E)$ would be null-homotopic. Hence, $g\circ f_{D(E)}$ factors through the quotient $D(E)\to D(E)/\partial D(E)$ up to homotopy. Note that $T:=D(E)/\partial D(E)$ is the Thom space of the disk bundle $D(E)$. Thus, we obtain a map $F:T\to S^{3}$. Composing the zero section $M\to D(E)$ with the quotient map $D(E)\to T$, we obtain an embedding $s:M\to T$. We observe $F\circ s=g\circ f_{D(E)}\circ s$ is homotopic to $g\circ f_{M}$. Thus, $F\circ s:M\to S^{3}$ is degree $1$, as well.
Henceforth, we closely imitate the proof of [Bol06, Theorem 2.1]. Let
$\Phi:H^{k}(M;\mathbf{Z}_{2})\to H^{k+2}(D(E),\partial D(E);\mathbf{Z}_{2})\cong H^{k+2}(T;\mathbf{Z}_{2})$
denote the Thom isomorphism, defined by $\Phi(x)=q^{*}x\smile u$ where $u\in H^{2}(T;\mathbf{Z}_{2})$ denotes the Thom class of the disk bundle $q:D(E)\to M$. Let $\eta\in H^{3}(S^{3};\mathbf{Z}_{2})\cong\mathbf{Z}_{2}$ be the generator and let $a=F^{*}(\eta)\in H^{3}(T;\mathbf{Z}_{2})$. Since $\Phi$ is an isomorphism, there exists $z\in H^{1}(M;\mathbf{Z}_{2})$ such that $\Phi(z)=a$. Let $w_{2}\in H^{2}(M;\mathbf{Z}_{2})$ denote the second Stiefel-Whitney class of $q:D(E)\to M$. Using $s^{*}u=w_{2}$ and the fact that $F\circ s:M\to S^{3}$ is degree $1$, we obtain
$z\smile w_{2}=s^{*}q^{*}z\smile s^{*}u=s^{*}(q^{*}(z)\smile u)=s^{*}(\Phi(z))=s^{*}(a)=(F\circ s)^{*}(\eta)\neq 0$ (3.3)
Let $\mathrm{Sq}^{i}$ denote the Steenrod square operations. Since $|z|=1$, we get $\mathrm{Sq}^{1}(z)=\mathrm{Sq}^{2}(z)=0$. Using Cartan’s formula, naturality of Steenrod squares and $|u|=2$, we compute
$\mathrm{Sq}^{2}(a)=\mathrm{Sq}^{2}(\Phi(z))=\mathrm{Sq}^{2}(q^{*}z\smile u)=q^{*}z\smile\mathrm{Sq}^{2}(u)=q^{*}z\smile u\smile u$ (3.4)
From Thom-Wu formula, we have $\Phi(w_{2})=u\smile u$. Plugging this back in (3.4), we obtain
$\mathrm{Sq}^{2}(a)=q^{*}z\smile\Phi(w_{2})=q^{*}z\smile q^{*}w_{2}\smile u=q^{*}(z\smile w_{2})\smile u=\Phi(z\smile w_{2})$ (3.5)
From (3.3), $z\smile w_{2}\neq 0$. Since $\Phi$ is an isomorphism, combining this with (3.5) implies $\mathrm{Sq}^{2}(a)\neq 0$. On the other hand, by naturality of Steenrod squares, we have
$\mathrm{Sq}^{2}(a)=\mathrm{Sq}^{2}(F^{*}\eta)=F^{*}\mathrm{Sq}^{2}(\eta)=0,$
since $\mathrm{Sq}^{2}(\eta)\in H^{5}(S^{3};\mathbf{Z}_{2})$ vanishes as $H^{5}(S^{3};\mathbf{Z}_{2})\cong 0$. This gives us the required contradiction. Therefore, $f_{E}$ is not homotopic into the $2$-skeleton of $K(G,1)$, as desired. ∎
We now proceed to the proof of Theorem B. As a crucial ingredient, we shall use a result due to He [He25, Theorem 1.9].
@@
Let $M^{3}$ be a closed $3$-manifold such that $M$ is either enlargeable or $\dim_{\mathrm{MC}}(\widetilde{M})=3$, and let $p:E\to M$ be a circle bundle over $M$ such that the total space of $E$ admits a PSC metric. As a first step, we argue that it can be ensured without loss of generality that $(1)$ $M$ is orientable, $(2)$ the circle bundle $E\to M$ is fiberwise orientable.
@@
    If $M$ is enlargeable, then $M$ is orientable by definition. If $\dim_{\mathrm{MC}}(\widetilde{M})=3$ and $M$ is not orientable, we pass to the orientation double cover of $M$ and pull the bundle $E\to M$ back to this cover. As the macroscopic dimension of a manifold is a quantity that only depends on the universal cover of the manifold, the cover continues to have macroscopic dimension $3$. Moreover, the total space of the pullback of $E\to M$ is a double cover of the total space of $E$. Therefore, the total space of the pulled back bundle also admits a PSC metric. Therefore, without loss of generality, we may assume $M$ is orientable.